Crane Company manufactures deep-sea fishing rods, which it distributes internationally through a chain of wholesalers. The following data are taken from the budget prepared at the beginning of the year by Crane’s controller. The company applies overhead on the basis of machine hours.

Annual Budget May Budget
Variable manufacturing overhead $3,373,500 $305,500
Fixed manufacturing overhead $1,203,600 $100,300
Direct labor hours 53,880 4,490
Machine hours 259,500 23,500


During the month of May, Crane used 4,280 direct labor hours and 21,930 machine hours. The flexible budget for the month allowed 4,420 direct labor hours and 21,000 machine hours. Actual fixed manufacturing overhead incurred was $103,300; variable manufacturing overhead incurred was $283,490.


(a) Calculate the variable overhead spending and efficiency variances for May. (Round per unit value to 2 decimal places, e.g. 52.75 and final answers to 0 decimal places, e.g. 5,725. If variance is zero, select "Not Applicable" and enter 0 for the amounts.)

Variable overhead spending variance

$_____________ Favorable

Variable overhead efficiency variance

$______________    Unfavorable
Answer #1

a) Variable overhead spending variance = (Standard rate-actual rate)actual hour = (13*21930-283490) = 1600 F

b) Variable overhead efficiency variance = (Standard hour-actual hour)Standard rate = (21000-21930)*13 = 12090 U

> how I can calculate the stanrad rate?

nancy adly Wed, Jan 19, 2022 2:53 PM